# FACE: Better Understanding Large Language Model Capabilities by Dissociating Language and Cognition

### v1

@@ 3 Experiments
Evaluation data construction. As summarized in Table 2, we collect a total of 17 English benchmarks and modify the corresponding input-output pairs into a unified QA format, i.e. generative QA or multiple-choice QA. The reference answers of the benchmarks are directly used as the final answer , while the reference rationales ( and ) for the intermediate reasoning steps are constructed manually. Specifically, on the one hand, , i.e. the reference rationale for the first reasoning step is based on the metadata analysis about subset division of the employed benchmarks. For example, when we evaluate the grammaticality regarding negative polarity item (NPI) licensing, we manually write “The word ’any’, in their most common uses, are negative polarity items: they can only be used in an appropriate syntactic-semantic environment—to a first approximation, in the scope of negation.” as the reference rationale for the whole “any”-based NPI licensing subset, other subset will substitute the “any” with corresponding licensing contexts or trigger words, such as “ever” and “even”. On the other hand, , i.e. the reference rationale for the second reasoning step is built on the instance-wise annotations of human evaluation publicly released by the authors of corresponding benchmarks, which annotates necessary explanations as well as final answer for a given question. Although this will leave few benchmarks available and lead to a limited number of evaluation data, it provides relatively reliable references and especially enables re-producible evaluation.
@@
Examined models. As summarized in Table 3, the examined LLMs can be categorized into publicly available open-source models and proprietary ones whose responses are provided through private APIs. On the one hand, open-source models include three backbone models, i.e. T5 Raffel et al. (2020), LLaMA Touvron et al. (2023a) and LLaMA 2 Touvron et al. (2023b), which are pre-trained on large scale corpus and not applied to any fine-tuning. Initialized with T5, Flan-T5 Longpre et al. (2023) and Flan-Alpaca Chia et al. (2023) are instruction-tuned on Flan V2 Longpre et al. (2023) and Alpaca Taori et al. (2023), respectively. Built on LLaMA, Alpaca Taori et al. (2023) and Vicuna Chiang et al. (2023) are instruction-tuned with responses generated by GPT-3.5, while TÜLU 1 Wang et al. (2023d) are instruction-tuned with a mixture of both manually curated and distilled dataset. Based on LLaMA 2, LLaMA 2-Chat Touvron et al. (2023b) is firstly instruction-tuned with high-quality collected annotations, and then aligned with human preferences for the chat use case. Besides, to perform a fair comparison w.r.t instruction-tuning dataset, we also evaluate LLaMA checkpoints fine-tuned on other datasets, such as Flan V2 Longpre et al. (2023) (human-written), Alpaca (model-generated) Taori et al. (2023), ShareGPT 111https://sharegpt.com/ (user prompt with model response). On the other hand, proprietary models consist of OpenAI’s GPT-3.5 (gpt-3.5-turbo) OpenAI (2022), OpenAI’s InstructGPT (gpt-3.5-turbo-instruct) Ouyang et al. (2022) and Google’s Bard Google (2023). Please refer to Appendix A for checkpoint details.
@@ 3.1 Main results
The difference in problem-solving performance between open-source and proprietary models is significantly greater than the difference in their crystallized performance. On the one hand, in terms of problem-solving performance (), as shown in Table 4, open-source models usually underperform the proprietary ones across various capabilities, especially cognition-related ones, such as world modeling and social modeling. For example, one of the most competitive open-source model Alpaca achieves a 49.42 accuracy in the world modeling dimension, while GPT-3.5 produces a performance of 67.68, excelling Alpaca by a substantial margin (around 36%). A similar conclusion can also be drawn from the other dimensions, such as the best problem-solving performance of proprietary models exceeds that of open-source models by about 16%, 25%, and 40% in linguistic knowledge, formal knowledge, and social modeling, respectively. On the other hand, in terms of crystallized performance (), there is a rather smaller gap between open-source and proprietary models compared to problem-solving accuracy. For example, the maximal difference of in the world modeling among all the examined models is about 9%, i.e. GPT-3.5’s 81.53 vs. T5’s 74.61, while their difference of is GPT-3.5’s 67.68 vs. T5’s 26.53. This kind of inconsistency between and can also be observed in other capability dimensions, potentially showing that either pre-training or supervised fine-tuning of LLMs can encode sufficient knowledge into the model, but the final task performance does not just depend on the amount or quality of knowledge.
@@
Language-related capabilities show a relatively weak correlation with cognitive capabilities. Figure 2 presents the correlation results between different capabilities. Despite all showing strong correlations (Pearson’s > 0.7 Krippendorff (2004)), both the language-related and cognition-related capabilities exhibit stronger intra-dimension correlation (e.g. world modeling vs.social modeling) when compared to inter-dimension correlation (e.g. world modeling vs. formal knowledge). This indicates that excellence in language processing does not necessarily equate to a similar level of cognitive capability. This result can also be observed from Table 4. For example, the model LLaMA 2-Chat achieves better results in terms of problem-solving performance of linguistics knowledge than Alpaca, i.e. LLaMA 2-Chat’s 74.46 vs. Alpaca’s 68.39, but it fails to surpass Alpaca in world modeling (44.46 vs. 49.42) and social modeling (29.59 vs. 32.96). This pattern also applies to other models, such as Bard vs. InstructGPT, further demonstrating the reasonability of disassociating language-related capabilities and cognition-related ones to better evaluate LLMs.
@@
A possible reason behind it could be that dedicated structures of the model or subsets of parameters are highly correlated with language, whereas others serve as cognition, and they are optimized at different training stages and function as different mechanisms during inference, which has been verified by recent studies on knowledge locating and editing of LLMs. For example, Zhao et al. (2023) and Chen et al. (2023) respectively found core linguistic regions and language-independent knowledge neurons in LLMs, and Dai et al. (2022); Meng et al. (2022) proposed to explicitly edit knowledge neurons to improve performance, potentially providing a feasible approach to strengthening LLMs capability without fine-tuning, i.e. editing their parameters or architectures directly.
@@
The crystallized step impacts problem-solving more than the fluid step. Figure 3 illustrates the relationship between problem-solving performance () and sum of crystallized () and fluid () performance. Both and make a difference to the final , showing that problem-solving not only depends on the amount or quality of stored knowledge but also is reflective of the effectiveness of knowledge utilization. For example, LLaMA 2 underperforms LLaMA 2-Chat in terms of across various capabilities. When taking a closer look at the intermediate results, we can observe that both the models do well in crystallized step, but LLaMA 2 shows an worse result in fluid step, leading the worse problem-solving performance than LLaMA 2-Chat. Besides, all of the open-source models exhibit relatively poor fluid performance w.r.t. GPT-3.5, especially those that are pre-trained but not instruction-tuned, such as T5 and LLaMA 2. This implies a solution to improve problem-solving performance, i.e. boosting the efficacy of knowledge utilization. Section 3.2 devise a knowledge-enhanced method to demonstrate this solution.
@@
Both the model size and the quality of fine-tuning dataset affect the capabilities of LLMs. Firstly, backbone models play a critical role in building superior models. For example, as shown in Table 4, fine-tuned on the same dataset, LLaMA-based Alpaca performs better than T5-based Flan-Alpaca, and the larger scale proprietary models show a greater advantage over other open-source models. In addition, scaling open-source models does improve both language and cognitive capability. As shown in Figure 4, problem-solving performance across various capabilities increases as the model size increases, and TÜLU-65B achieves the best performance. In particular, the level of formal knowledge of TÜLU-65B are close to that of GPT-3.5. Last, but not least, there is no significant performance difference among various open-source instruction-tuning datasets whether it is comprised of human-written instruction or not. As illustrated in Figure 5, there is not a single best instruction tuning dataset across all tasks, indicating different datasets bring different benefits to LLMs’ capabilities. This finding is consistent with the recent success of a mixture of instruction-tuning datasets or expert LLMs Jiang et al. (2024); Xia et al. (2024).
@@ 3.2 Boosting LLMs with Injected Knowledge
Based on the above analysis showcasing the limitations of the crystallized performance of existing LLMs, as illustrated in Figure 6, we propose to a knowledge-enhanced approach. Specifically, for each given instance with question and answer, the first baseline, denoted as +, append the first reference rationale, i.e. , to the input question with string concatenation. Then the augmented input is fed into the model with the same instruction as the examined model. Note that we also remove the first triplet of [thought], [action], [answer] in the input demonstrations for the + baseline because we have provided the corresponding reference rationale. As a comparison, following a similar procedure, we also construct another baseline by incorporating both and into the model, denoted as ++.
@@ 4 Related Works
Evaluation of LLMs. LLMs are initially assessed on various understanding Wang et al. (2018, 2019) and generation tasks Pilault et al. (2020); Thompson and Post (2020). With the increasing emphasis on the trustworthiness of models, dedicated benchmarks are proposed to evaluate robustness Yang et al. (2022); Wang et al. (2023b), hallucination Li et al. (2023), bias Zhong et al. (2023), and generalizability Wang et al. (2023a). In a more unified form, recent-emerged works aim to evaluate LLMs on holistic benchmark Mialon et al. (2023); Rein et al. (2023). Chia et al. (2023) conduct evaluation from problem-solving, writing, and human alignments, while Ye et al. (2023) annotates a single instance with a set of skills, including logical thinking, background knowledge, problem handling, and user alignment. Although they provide fine-grained analysis of LLMs’ capability, they quantify each capability only using the performance metric, hence suffering from issues of interpretability.
Cognition-inspired intelligence evaluation. How to define and evaluate intelligence is widely investigated by both cognitive science and AI benchmark design Cattell (1963); Rogers et al. (2023). Taking MRC as the target, Chollet (2019) describes intelligence as skill-acquisition efficiency, while Sugawara et al. (2020) and Ray Choudhury et al. (2022) respectively propose to benchmark MRC based on reasoning skills and steps a system would be “reading slowly”. As for the evaluation of LLMs, Mahowald et al. (2023) summarize lots of neuroscience evidence for dissociating language and thought, largely motivating the design of FACE.
@@ 5 Conclusion
We present FACE, defining a fine-grained capability evaluation framework for LLMs by dissociating language and cognition. FACE decomposes each capability into sub-steps to assess performance of knowledge recalling and utilization as well as problem-solving. FACE reveals the inherent limitations of existing LLMs in knowledge utilization and provides a knowledge-enhanced remedy for it. Empirical results demonstrate its effectiveness.
@@
Our work proposes a fine-grained and cognition-grounded capability evaluation framework for LLMs, namely FACE, which is based on the dissociated relationship of language and cognition, and evaluating the intermediate reasoning steps of LLMs. The limitations are two-fold, including data quality and domain generalizability.
On the other hand, motivated by a variety of empirical evidence from both neuroscience and probing experiments of LLMs, we formulate FACE as four capability dimensions, then re-formulate instances from multiple existing benchmarks and conduct stepwise evaluation. Although we try to ensure that the employed datasets are as consistent and targeted with the defined dimensions as possible, the kind of evaluation data construction might not reflect the required skills accurately. For example, an instance can cover more than one language or cognition capabilities As discussed in Section 3.1, different capabilities are correlated to each other to some extent, Besides, the reference rationales, i.e. the gold standard of the intermediate reasoning step, are based on the human annotations from original benchmarks, leading to the inconsistency of reference answers and a limited number of available data, which might bias the evaluation results. One remedy to these incidental issues could be building a new holistic benchmark with fine-grained annotations following our proposed schema. We regard it as our future work and deem designing a new annotation specification a promising direction.
On the other hand, our FACE only examined LLMs on general domains and English input, ignoring the domain-specific and multilingual application. In particular, the reasoning process of LLMs may expose social bias encoded in these models, such as race and gender Lucy and Bamman (2021). Therefore, additional evaluation protocols considering potential risks to user safety are left for our future work.
@@
We introduce FACE, a fine-grained and cognition-grounded capability evaluation framework for LLMs, and conduct evaluation experiments on publicly available datasets which are widely used in related research. Although LLMs have the potential to cause harm at the individual and societal levels Gonen and Goldberg (2019), our FACE aims to provide a deep understanding of the capabilities and limitations of LLMs, potentially making the risks from the LLMs more predictable.

### v2

@@ 3 Experiments
Evaluation data construction. In Table 2, we present a collection of 17 widely adopted English benchmarks and modify the corresponding input-output into a unified QA format, i.e. generative QA or multiple-choice QA. The choice to utilize these benchmarks is rooted in considerations of data quality, availability, and specificity of focus; hence, some widely recognized benchmarks may not be included for these reasons. For example, PIQA Bisk et al. (2020) focuses on physical commonsense, which, while valuable, represents only a single facet of commonsense reasoning. In contrast, MMLU Hendrycks et al. (2020) encompasses a broad spectrum of subjects, but requires both commonsense and contextual understanding, which might not align with our goal of ensuring a broad range of capabilities while minimizing the coupling between abilities during evaluation.
The reference answers of the benchmarks are directly used as the final answer , while the reference rationales ( and ) for the intermediate reasoning steps are constructed automatically. Specifically, on the one hand, , i.e. the reference rationale for the first reasoning step, is based on the rationale templates and the gold labels of the employed benchmarks. For example, when evaluating the grammaticality regarding negative polarity item (NPI) licensing, the rationale template for could be “The word is a negative polarity item: it can only be used in the scope of negation.”. The blank is then filled with gold labels (licensing contexts or trigger words), such as “any”, “ever”, and “even”, to build the final for corresponding NPI licensing samples. On the other hand, , i.e. the reference rationale for the second reasoning step is built on the instance-wise annotations of human evaluation publicly released by the authors of corresponding benchmarks, which annotates necessary explanations as well as final answer for a given question. Although this will leave few benchmarks available and lead to a limited number of evaluation data, it provides relatively reliable references and especially enables reproducible evaluation.
@@
| LLaMA 3-Instruct | 8B | 15.0T | IT+RLHF |
| LLaMA 3.1-Instruct | 8B | 16.4T | IT+RLHF |
@@
Examined models. As summarized in Table 3, the examined LLMs can be categorized into publicly available open-source models and proprietary ones whose responses are provided through private APIs. Open-source models include three backbone models, i.e. T5 Raffel et al. (2020), LLaMA Touvron et al. (2023a) and LLaMA 2 Touvron et al. (2023b), which are pre-trained on large scale corpus and not applied to any fine-tuning. Initialized with T5, Flan-T5 Longpre et al. (2023) and Flan-Alpaca Chia et al. (2023) are instruction-tuned on Flan V2 Longpre et al. (2023) and Alpaca Taori et al. (2023), respectively. Built on LLaMA, Alpaca Taori et al. (2023) and Vicuna Chiang et al. (2023) are instruction-tuned with responses generated by GPT-3.5, while TÜLU 1 Wang et al. (2023d) are instruction-tuned with a mixture of both manually curated and distilled dataset. Based on LLaMA 2, LLaMA 2-Chat Touvron et al. (2023b) is firstly instruction-tuned with high-quality collected annotations, and then aligned with human preferences for the chat use case. LLaMA 3-Instruct and LLaMA 3.1-Instruct respectively fine-tune LLaMA 3 and LLaMA 3.1 to better understand and follow human instructions. Besides, to perform a fair comparison w.r.t instruction-tuning dataset, we also evaluate LLaMA checkpoints fine-tuned on other datasets, such as Flan V2 Longpre et al. (2023) (human-written), Alpaca (model-generated) Taori et al. (2023), ShareGPT (user prompt with model response).
Proprietary models consist of OpenAI’s GPT-3.5 (gpt-3.5-turbo) OpenAI (2022), InstructGPT (gpt-3.5-turbo-instruct) Ouyang et al. (2022), GPT-4 (gpt-4-turbo) Achiam et al. (2023), and Google’s Bard Google (2023) (also known as Gemini Team et al. (2023)).
@@ 3.1 Main results
The difference in problem-solving performance is significantly greater than the difference in crystallized performance. On the one hand, in terms of problem-solving performance (), as shown in Table 4, open-source models usually underperform the proprietary ones across various capabilities, especially cognition-related ones, such as world modeling and social modeling. For example, the most competitive open-source model LLaMA 3.1-Instruct achieves a 50.39 accuracy in the world modeling dimension, while GPT-4 produces a performance of 69.43, excelling Alpaca by a substantial margin (around 40%). A similar conclusion can also be drawn from the other dimensions, such as the best problem-solving performance of proprietary models exceeds that of open-source models by about 20%, 32%, and 40% in linguistic knowledge, formal knowledge, and social modeling, respectively. On the other hand, in terms of crystallized performance (), there is a rather smaller gap between open-source and proprietary models compared to problem-solving accuracy. For example, the maximal difference of in the world modeling among all the examined models is about 9%, i.e. GPT-4’s 81.86 vs. T5’s 74.61, while their difference of is GPT-4’s 69.43 vs. T5’s 26.53. This inconsistency between and can also be observed in other capability dimensions, potentially showing that either pre-training or fine-tuning of LLMs can encode sufficient knowledge into the model, but the final task performance does not just depend on the amount or quality of knowledge.
Notably, the crystallized performance (s1) of LLaMA 3-Instruct and LLaMA 3.1-Instruct has significantly improved compared to LLaMA 2-Chat. For example, in social modeling, LLaMA 3-Instruct scored 76.99, and LLaMA 3.1-Instruct scored 77.78, compared to LLaMA 2-Chat’s 71.06. This suggests that these models better encode knowledge relevant to higher-level cognitive tasks, likely due to training and fine-tuning LLaMA 3 on higher-quality and larger-scale data compared to the LLaMA 2 model.
@@
Linguistic capabilities show a relatively weak correlation with cognitive capabilities. Figure 2 presents the correlation results between different capabilities. Both the language-related and cognition-related capabilities exhibit stronger (Pearson’s > 0.7 Krippendorff (2004)) intra-dimension correlation (e.g. world modeling vs. social modeling) when compared to inter-dimension correlation (e.g. world modeling vs. linguistic knowledge). This indicates that excellence in language processing does not necessarily equate to a similar level of cognitive capability. These results can also be observed from direct prompt evaluation, which assesses problem-solving performance () without prompting intermediate reasoning steps, further demonstrating the rationality of dissociating language and cognition.
@@
A possible reason behind it could be that dedicated structures of the model or subsets of parameters are highly correlated with language Zhang et al. (2024b); Tang et al. (2024), whereas others serve as cognition Chen et al. (2023), and they are optimized at different training stages and function as different mechanisms during inference, which has been verified by recent studies on knowledge locating and editing of LLMs Dai et al. (2022); Meng et al. (2022); Zhang et al. (2024a).
@@
The crystallized step impacts problem-solving more than the fluid step. Figure 3 illustrates the relationship between problem-solving performance () and sum of crystallized () and fluid () performance. Both and make a difference to the final , showing that problem-solving not only depends on the amount or quality of stored knowledge but also is reflective of the effectiveness of knowledge utilization. For example, LLaMA 2 underperforms LLaMA 2-Chat in terms of across various capabilities. When taking a closer look at the intermediate results, we can observe that both the models do well in crystallized step, but LLaMA 2 shows an worse result in fluid step, leading the worse problem-solving performance than LLaMA 2-Chat. Besides, all of the open-source models exhibit relatively poor fluid performance w.r.t. GPT-3.5, especially those that are pre-trained but not instruction-tuned, such as T5 and LLaMA 2. This implies a solution to improve problem-solving performance, i.e. boosting the efficacy of knowledge utilization. Section 3.2 presents a knowledge-enhanced method to demonstrate this solution.
@@
Both the model size and the quality of fine-tuning dataset affect the capabilities of LLMs. Firstly, backbone models play a critical role in building superior models. For example, as shown in Table 4, fine-tuned on the same dataset, LLaMA-based Alpaca performs better than T5-based Flan-Alpaca, and the larger scale proprietary models show a greater advantage over other open-source models. In addition, scaling open-source models does improve both language and cognitive capability. As shown in Figure 4, problem-solving performance across various capabilities increases as the model size increases, and 65B achieves the best performance. In particular, the level of formal knowledge of 65B is close to that of GPT-3.5. Last, but not least, there is no significant performance difference among various open-source instruction-tuning datasets whether it is comprised of human-written instruction or not. As illustrated in Figure 5, there is not a single best instruction tuning dataset across all tasks, indicating different datasets bring different benefits to LLMs’ capabilities. This finding is consistent with the recent success of a mixture of instruction-tuning datasets or expert LLMs Jiang et al. (2024); Xia et al. (2024).
@@ 3.2 Boosting LLMs with Injected Knowledge
Based on the above analysis showcasing the limitations of the crystallized performance of existing LLMs, as illustrated in Figure 6, we propose a knowledge-enhanced approach. Specifically, for each given instance with a question and answer, the first baseline, denoted as +, append the first reference rationale, i.e. , to the input question with string concatenation. Then the augmented input is fed into the model with the same instruction as the examined model. Note that we also remove the first triplet of [thought], [action], [answer] in the input demonstrations for the + baseline because we have provided the corresponding reference rationale. As a comparison, following a similar procedure, we also construct another baseline by incorporating both and into the model, denoted as ++.
@@ 4 Related Works
Evaluation of LLMs. LLMs are initially assessed on various understanding Wang et al. (2018, 2019) and generation tasks Pilault et al. (2020); Thompson and Post (2020). With the increasing emphasis on the trustworthiness of models, dedicated benchmarks are proposed to evaluate robustness Yang et al. (2022); Wang et al. (2023b), hallucination Li et al. (2023); Belyi et al. (2024), and generalizability Wang et al. (2023a). Recent-emerged works evaluate LLMs using another evaluator LLM Kim et al. (2024a, b) or on a holistic benchmark Mialon et al. (2023); Rein et al. (2023). Chia et al. (2023) conduct evaluation from problem-solving, writing, and human alignments, while Ye et al. (2024) annotates a single instance with a set of skills, including logical thinking, background knowledge, problem handling, and user alignment. Although they provide fine-grained analysis of LLMs’ capability, they suffer from the limited number of testing instances, e.g. 1,700 of Ye et al. (2024) and overlook the language-related low-level capabilities.
Cognition-inspired intelligence evaluation. How to define and evaluate intelligence is widely investigated by both cognitive science and AI benchmarking Cattell (1963); Rogers et al. (2023). In the MRC evaluation, Chollet (2019) describes intelligence as skill-acquisition efficiency, while Sugawara et al. (2020), Wang et al. (2022a) and Ray Choudhury et al. (2022) propose to benchmark MRC through reasoning skills and steps a system would be “reading slowly”. As for the LLMs, Mahowald et al. (2023) summarize extensive neuroscience evidence of human language and propose a conceptual framework with formal and functional competencies, which largely motivated the design of FAC2E. Compared to Mahowald et al. (2023), FAC2E introduces more concrete NLP tasks with a more comprehensive capability system and leverages stepwise evaluation to improve the accuracy of assessments.
@@ 5 Conclusion
We present FAC2E, defining a fine-grained capability evaluation framework for LLMs, which decomposes each capability into sub-steps to assess performance of knowledge recalling and utilization as well as problem-solving. FAC2E reveals the limitations of existing LLMs in knowledge utilization and provides a knowledge-enhanced remedy for it. Empirical results demonstrate its effectiveness.
@@
Our work proposes a fine-grained and cognition-grounded capability evaluation framework for LLMs, namely FAC2E, which is based on the dissociated relationship of language and cognition, and evaluating the intermediate reasoning steps of LLMs. The limitations are two-fold, including data quality and domain generalizability.
On the other hand, motivated by a variety of empirical evidence from both neuroscience and probing experiments of LLMs, we formulate FAC2E as four capability dimensions, then re-formulate instances from multiple existing benchmarks and conduct stepwise evaluation. Although we try to ensure that the employed datasets are as consistent and targeted with the defined dimensions as possible, the kind of evaluation data construction might not reflect the required skills accurately. For example, an instance can cover more than one language or cognition capabilities As discussed in Section 3.1, different capabilities are correlated to each other to some extent, Besides, the reference rationales, i.e. the gold standard of the intermediate reasoning step, are based on the human annotations from original benchmarks, leading to the inconsistency of reference answers and a limited number of available data, which might bias the evaluation results. One remedy to these incidental issues could be building a new holistic benchmark with fine-grained annotations following our proposed schema. We regard it as our future work and deem designing a new annotation specification a promising direction.
On the other hand, our FAC2E only examined LLMs on general domains and English input, ignoring the domain-specific and multilingual application. In particular, the reasoning process of LLMs may expose social bias encoded in these models, such as race and gender Lucy and Bamman (2021). Therefore, additional evaluation protocols considering potential risks to user safety are left for our future work.
@@
We introduce FAC2E, a fine-grained and cognition-grounded capability evaluation framework for LLMs, and conduct evaluation experiments on publicly available datasets which are widely used in related research. Although LLMs have the potential to cause harm at the individual and societal levels Gonen and Goldberg (2019), our FAC2E aims to provide a deep understanding of the capabilities and limitations of LLMs, potentially making the risks from the LLMs more predictable.
Acknowledgements
This work is supported by the Canada CIFAR AI Chair Program and the Canada NSERC Discovery Grant (RGPIN-2021-03115).